Circular No. 7176

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

NOVA VELORUM 1999 Independent visual discoveries of a bright nova have been made by

Peter Williams, Heathcote, New South Wales (via S. Lee and A. Pearce), at m_{v} = 3.1 on May 22.396 UT and by Alan C. Gilmore, Mount John University Observatory, on May 22.451. Lee provides the following position from Anglo-Australian Telescope (AAT) encoders: α = 10^{h}44^{m}49^{s}.5, δ = −52^{o}25′35′′ (equinox 2000.0). Lee reports that a low-dispersion spectrogram taken by Lewis, Colless, Cannon, Bridges, and himself with the AAT (+ fiber spectrograph on 2dF; 1-s exposure, mirror stopped down to 2.5 m) shows Hα and Hβ in narrow emission with a strong P-Cyg profile. An échelle spectrogram (10-min exposure centered on May 22.500) of the apparent nova by K. R. Pollard and J. A. McSaveney with the Mt. John 1-m telescope shows Hα with blueshifted absorption and slightly redshifted emission. Gilmore provides the following photometry obtained with the Mount John 0.6-m reflector (comparison star HR 4167 = p Vel, with assumed V = 3.84, U − B = +0.07, B − V = +0.30, V − R = +0.20, V − I = +0.30): May 22.488, V = 2.88, U − B = −0.27, B − V = +0.24, V − R = +0.16, V − I = +0.28; 22.574, 2.85, −0.27, +0.28, +0.19, +0.27; 22.599, 2.84, −0.25, +0.29, +0.23, +0.33; 22.635, 2.80, −0.26, +0.31, +0.23, +0.33. P. M. Kilmartin reports that a 1-s CCD (over)exposure on May 22.50 with the 0.6-m Boller & Chivens Cassegrain reflector at Mount John yields the following position end figures: 48^{s}.1, 32′′ (four stars, mean residual 2′′.68 in α, 0′′.43 in δ). Pearce provides the following visual magnitude estimates: May 22.42–22.52, 3.0; 22.54–22.58, 2.9. Pearce notes that there appears to be nothing obvious on the Digital Sky Survey image at the nova’s position. P. Caldwell, Geelong, Victoria, reports m_{v} ∼ 2.8 (close to or slightly fainter than the brightness of μ Vel) during May 22.50–22.53.

COMET C/1999 K4 (LINEAR) An apparently asteroidal object found by LINEAR (discovery obser-

vation below; further observations and orbit on MPEC 1999-K24), and posted on The NEO Confirmation Page, has been reported as cometary by M. Hicks (Table Mountain; faint coma of diameter ∼ 5′′ on May 21) and by C. W. Hergenrother and A. E. Gleason (Catalina 1.5-m reflector; highly condensed coma with a faint 10′′–15′′ tail in p.a. 170^{o} on May 22).

1999 UT  α_{2000}  δ_{2000}  m_{2}
May 17.32636  16^{h}05^{m}48.^{s}69  +5^{o}45′19′′.7  18.7
1999 May 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green